# Identifying treatment response subgroups in observational time-to-event data

### v1

@@ A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in Eq. (3).
@@ A.2 Upper-Bound Simplification
with $A$, being the treatment assignment. The following derives this result:
$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$ (Under Assumption 3)
@@ B.1 Data generation
We consider a synthetic population of $N=30,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to [34]. The following describes our generation process:
@@
$X_{[3:10]}$ $\sim\text{MVN}(0,I^{8})$
with MVN denoting a multivariate normal distribution, and $I^{n}$, the identity covariance matrix of dimension $n$.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications. After drawing the treatment assignment probability $P$, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational". "Randomised" consists of a Bernoulli draw using the realisation of $P$. "Observational" reflects an assignment dependent upon the observed covariates.
$P$ $\sim\text{Uniform}(0.25,0.75)$
$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(p)$
$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times p)$
with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$. In our experiment, we chose $\Phi(x)=x[1]^{2}+x[2]^{2}$.
@@
B.2 Optimisation grid search
@@
We adopted a 10-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[\![25,50,100]\!]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $250$ or $500$.
@@
Appendix C Case Study: SEER dataset
@@
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^9] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following [25, 10, 20], we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 tumour characteristics measured at diagnosis.
@@
From the selected population, our aim is to identify heterogeneous responses to adjuvant treatment. The first challenge is the selection of the number of groups to use ($K$). We advise to follow medical actionability and consider the change in treatment effects and size of the subgroups when increasing this parameter. In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$. This means selecting the number of subgroups after which one observes limited likelihood improvement as used in [19]. Using Figure 5, the negative log-likelihood presents an elbow for $K=2$.
@@
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist [24], making this setting more likely to meet the positivity assumption (Assumption 6), necessary to study causality in observational data.
@@
(b) Cox Mixture of Heterogenous Effects
@@
Figure 6 presents the identified treatment effect subgroups when using CNSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning. As shown by Figure 5, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
@@
Figure 7: Causal Neural Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
Table 2 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) [42]. Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
@@
Table 2: Causal Neural Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy. However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options.

### v3

@@ A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in equation 3.
@@ A.2 Upper-Bound Simplification
with $A$, being the treatment assignment. Under assumption 3, expressing this quality corresponds to:
$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$
@@ B.1 Data generation
We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al. (2022a). The following describes our generation process:
@@
$Y$ $\sim\text{Mult}\left(1,\left[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}\right]\right)$
$X_{[3:10]}\mid Y=k$ $\sim\text{MVN}(c^{\prime}_{k},I^{8})$
with MVN denoting a multivariate normal distribution, $c^{\prime}_{k}$ random cluster centres, and $I^{n}$, the identity covariance matrix of dimension $n$. Note that we introduce $Y$ to present a covariates structure that is independent of the treatment responses.
@@
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications. Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational". "Randomised" consists of a Bernoulli draw using the realisation of $P$. "Observational" reflects an assignment dependent upon the observed covariates.
$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(0.5)$
$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$. In our experiment, we chose $\Phi(x)=\sum x^{2}$.
@@
B.2 Training and hyperparameter optimisation
Training.
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2. All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
Hyperparameter optimisation.
@@ B.2 Training and hyperparameter optimisation
We adopted a 100-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
B.3 Sensitivity
In Section 3, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$. This section explores how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters. This analysis examines the risk associated with misspecifying the number of clusters.
Figure 5 illustrates the cross-validated subgroup treatment effects when trained with $K=2,3$ and $4$ clusters. Uncertainty increases with misspecified models due to the instability of the identified subgroups.
(a) $K=2$
(b) $K=3$ (as in the main text)
(c) $K=4$
Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
B.4 Alternative data generations
In this section, we explore alternative data generations to demonstrate the robustness of the proposed strategy.
B.4.1 Unequal cluster size
Section B.1 presents a population equally distributed over the different clusters. In medical settings, the underlying subgroups may differ in size. This section explores an alternative scenario in which the population is distributed over the three clusters: 62.5%, 25% and 12.5%.
Table 2 summarises the performance in this simulation, echoing the main text’s conclusions. The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects..
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.637 (0.068) 0.064 (0.013) 0.021 (0.009) 0.062 (0.010) CSC Unadjusted 0.651 (0.049) 0.055 (0.009) 0.017 (0.010) 0.067 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.780 (0.086) 0.064 (0.027) 0.040 (0.011) 0.069 (0.010) CSC Unadjusted 0.616 (0.146) 0.072 (0.020) 0.039 (0.028) 0.117 (0.033) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
Table 2: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
B.4.2 Number of points
Section B.1 describes a data generation with $3,000$ patients. This section presents results when there are $300$ and $30,000$ patients. A smaller number of patients may result in a larger impact of treatment non-randomisation but also impact the neural network capacity to identify underlying subgroups of treatment effect. A larger number makes non-randomisation less of a concern.
Table 3 presents the performance with these different population sizes under an observational treatment assignment. A first observation is that all methodologies present better performance with larger number of points. Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimate with a smaller population. This difference decreases when $N=30,000$ with the virtual twins approach presenting the best recovery of the treatment effects. However, throughout the different settings, the proposed CSC presents the best Rand-Index indicating a good recovery of the underlying structure.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.504 (0.139) 0.131 (0.090) 0.085 (0.033) 0.091 (0.062) CSC Unadjusted 0.449 (0.106) 0.202 (0.049) 0.162 (0.055) 0.125 (0.088) CMHE ($K=3$) 0.126 (0.084) 0.340 (0.050) 0.135 (0.043) 0.166 (0.050) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.087 (0.118) 0.284 (0.056) 0.639 (0.939) 0.747 (1.014) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.774 (0.015) 0.018 (0.010) 0.012 (0.003) 0.030 (0.009) CSC Unadjusted 0.737 (0.022) 0.033 (0.006) 0.010 (0.002) 0.039 (0.008) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
Table 3: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
B.4.3 Number of clusters
Section B.1 describes a data generation with $K=3$ clusters. This section explores the behaviour of the different methodologies with an increased number of clusters $K=5$. Following the same data generation, we introduces two additional clusters with centres [-3, 3] and [4, 4].
Table 4 presents the different models’ performance in this setting, echoing the same results as the main text. The proposed method best recovers the underlying clustering structure as shown by the Rand-Index. CMHE suffers from its inherent association between treatment and survival. The step-wise approach is unable to identify the clusters of interest if the data presents a structure that is not fully aligned with the outcome of interest.
This scenario further validates our findings, demonstrating the methodology’s capacity to generalise beyond the scenario presented in the main text.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ Randomised CSC 0.475 (0.038) 0.073 (0.030) 0.059 (0.023) 0.041 (0.015) 0.099 (0.024) 0.017 (0.003) CSC Unadjusted 0.459 (0.037) 0.095 (0.055) 0.054 (0.027) 0.046 (0.011) 0.076 (0.020) 0.019 (0.005) CMHE ($K=5$) 0.306 (0.016) 0.091 (0.036) 0.104 (0.012) 0.015 (0.002) 0.093 (0.011) 0.232 (0.013) CMHE ($L=5$) 0.256 (0.018) 0.135 (0.020) 0.112 (0.008) 0.020 (0.003) 0.221 (0.009) 0.349 (0.006) CMHE ($K=L$) 0.372 (0.018) 0.101 (0.023) 0.181 (0.013) 0.072 (0.003) 0.119 (0.010) 0.323 (0.010) KMeans + TE 0.156 (0.012) 0.075 (0.009) 0.184 (0.016) 0.072 (0.012) 0.147 (0.014) 0.039 (0.010) Virtual Twins 0.284 (0.081) 0.034 (0.029) 0.066 (0.026) 0.029 (0.012) 0.082 (0.023) 0.036 (0.010) Observational CSC 0.446 (0.016) 0.119 (0.072) 0.087 (0.018) 0.026 (0.008) 0.097 (0.023) 0.027 (0.016) CSC Unadjusted 0.444 (0.048) 0.077 (0.046) 0.065 (0.026) 0.045 (0.020) 0.118 (0.026) 0.034 (0.022) CMHE ($K=5$) 0.303 (0.010) 0.095 (0.030) 0.112 (0.017) 0.018 (0.005) 0.094 (0.013) 0.229 (0.010) CMHE ($L=5$) 0.217 (0.026) 1.341 (2.717) 0.119 (0.014) 0.022 (0.005) 1.432 (2.720) 1.561 (2.709) CMHE ($K=L$) 0.371 (0.029) 0.105 (0.009) 0.208 (0.043) 0.076 (0.006) 0.186 (0.025) 0.325 (0.004) KMeans + TE 0.162 (0.011) 0.058 (0.005) 0.210 (0.031) 0.084 (0.018) 0.122 (0.032) 0.032 (0.006) Virtual Twins 0.320 (0.084) 0.062 (0.038) 0.026 (0.008) 0.044 (0.021) 0.077 (0.040) 0.038 (0.018)
Table 4: Cross-validated performance (with standard deviation in parenthesis) when $K=5$.
B.4.4 Impact of treatment rate
Section B.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
Table 5 presents the performance under these different treatment rates in an observational setting. These results highlight CSC’s capacity to identify the underlying subgroups with the highest Rand-Index and one of the best cluster treatment effect recovery in these settings.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.790 (0.036) 0.036 (0.009) 0.034 (0.009) 0.018 (0.006) CSC Unadjusted 0.849 (0.033) 0.040 (0.005) 0.019 (0.005) 0.014 (0.004) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.149 (0.080) 0.188 (0.004) - 0.154 (0.005) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.808 (0.043) 0.035 (0.012) 0.027 (0.006) 0.019 (0.007) CSC Unadjusted 0.687 (0.048) 0.045 (0.011) 0.035 (0.016) 0.028 (0.006) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.010) 0.077 (0.005) 0.069 (0.004) CMHE ($L=3$) 0.318 (0.140) - - 0.132 (0.008) CMHE ($K=L$) 0.483 (0.104) 0.208 (0.013) - - KMeans + TE 0.002 (0.003) 0.195 (0.015) 0.079 (0.012) 0.148 (0.011) Virtual Twins 0.570 (0.013) 0.030 (0.005) 0.024 (0.013) 0.022 (0.005)
Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates.
B.4.5 Aligned covariates and treatment effect structure
Section B.1 presents a data generation with the last 8 covariates having a clustered structures independent of the treatment response. This section explores a setting where all clusters in the covariates are associated with different treatment responses of interest. Specifically, we sample the last dimensions from standard normal distributions instead of various clusters, i.e.,
$X_{[3-10]}\sim\text{MVN}(0,I^{8})$
Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach. When covariates’ clusters are aligned with the outcome of interest, this method recovers the clustering structure well, as shown by the best Rand-Index. Non-randomisation of treatment, however, worsens the treatment estimates due to the absence of correction in the estimation of treatment responses.
Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses. This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across settings.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.487 (0.052) 0.027 (0.013) 0.047 (0.020) 0.030 (0.009) CSC Unadjusted 0.522 (0.041) 0.025 (0.004) 0.027 (0.018) 0.034 (0.013) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.289 (0.078) 0.063 (0.009) 0.050 (0.019) 0.146 (0.020) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.666 (0.146) 0.016 (0.007) 0.029 (0.013) 0.033 (0.010) CSC Unadjusted 0.613 (0.207) 0.024 (0.016) 0.039 (0.015) 0.030 (0.003) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.294 (0.123) 0.057 (0.005) 0.061 (0.007) 0.144 (0.007) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses. This setting aligns with the assumptions made by the step-wise approaches.
B.4.6 Linear treatment responses
As an alternative setting, we propose a setting with a linear treatment response, as assumed by CMHE Nagpal et al. (2023). To this end, we followed a similar process as described in Section B.1 at the exception of the treatment response. We draw the response under no treatment from a Gompertz, and the one under treatment with the same parameters shifted by the treatment effect. Formally, using the same notation
$B^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{0}_{z},A_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0},a_{z})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x)\times a_{z},s_{0}(\gamma_{z}^{0},x)\right)$
with $a_{z}$ the treatment effect for cluster $z$. In this experiments, we choose $a_{0},a_{1},a_{2}=(0.5,1,1.5)$.
Table 7 summarises performances in this scenario. While demonstrating a significant improvement of the CMHE alternatives, CSC presents competing performance on subgroup identification and treatment effect estimates.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.389 (0.065) 0.033 (0.015) 0.058 (0.035) 0.014 (0.006) CSC Unadjusted 0.411 (0.125) 0.029 (0.017) 0.062 (0.021) 0.017 (0.015) CMHE ($K=3$) 0.304 (0.033) 0.040 (0.003) 0.050 (0.003) 0.030 (0.003) CMHE ($L=3$) 0.211 (0.109) 0.051 (0.003) 0.106 (0.006) 0.028 (0.008) CMHE ($K=L$) 0.384 (0.060) 0.411 (0.797) - 0.435 (0.650) KMeans + TE 0.101 (0.139) 0.080 (0.011) 0.064 (0.028) 0.038 (0.023) Virtual Twins 0.265 (0.085) 0.025 (0.006) 0.018 (0.004) 0.064 (0.013) Observational CSC 0.311 (0.043) 0.029 (0.003) 0.074 (0.028) 0.027 (0.016) CSC Unadjusted 0.275 (0.168) 0.032 (0.005) 0.044 (0.044) 0.026 (0.014) CMHE ($K=3$) 0.360 (0.068) 0.038 (0.002) 0.050 (0.002) 0.035 (0.009) CMHE ($L=3$) 0.285 (0.111) 0.055 (0.006) 0.094 (0.007) 0.017 (0.009) CMHE ($K=L$) 0.382 (0.058) 0.052 (0.011) 0.128 (0.013) 0.028 (0.020) KMeans + TE 0.115 (0.154) 0.068 (0.020) 0.076 (0.031) 0.031 (0.008) Virtual Twins 0.186 (0.117) 0.048 (0.027) 0.041 (0.010) 0.039 (0.012)
Table 7: Cross-validated performance (with standard deviation in parenthesis) when considering a linear treatment response. This setting aligns with the assumption made by CMHE.
@@ B.4.6 Linear treatment responses
Appendix C Real-world case Study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
@@
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following Lee et al. (2018); Danks & Yau (2022); Jeanselme et al. (2023), we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  Danks & Yau (2022) for further description).
@@
From the selected population, our aim is to identify heterogeneous responses to adjuvant treatment. The first challenge is the selection of the number of groups to use ($K$). We advise to follow medical actionability and consider the change in treatment effects and size of the subgroups when increasing this parameter. In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$. Using Figure 6, the negative log-likelihood presents an elbow for $K=2$.
@@
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist Lazzari et al. (2023), making this setting more likely to meet the positivity assumption (assumption 6), necessary to study causality in observational data.
@@
(b) Cox Mixture of Heterogeneous Effects
@@
Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning. As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
@@
Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
Table 8 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) Royston & Parmar (2013). Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
@@
Table 8: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy. However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options. The quality of available covariates limits this analysis and only serves as an example to medical practitioners to identify subgroups of treatment responses from observational data.